# Lizard: An Efficient Linearization Framework for Large Language Models

### v2

@@ 3.1 First Stage: Approximating Softmax Attention for Unbounded Context
where $N$ is the number of attention layers. Overall, Lizard achieves an $O\bigl{(}L(w+m)d+Ld^{2}\bigr{)}$ time and space complexity. For generation, Lizard requires only $O((w+m)^{2}d+d^{2})$ time and space for every token.
3.2 Second Stage: Aligning Subquadratic Attention with Language Modeling
While the first stage focuses on minimizing the approximation error between the hybrid subquadratic attention and full softmax attention, it does not directly optimize for the final language modeling performance. To bridge the gap between architectural approximation and practical performance, we fine-tune the model on autoregressive language modeling objectives to align the approximated attention mechanism with downstream performance. The loss function used during this stage is the standard autoregressive language modeling objective: $\mathcal{L}_{\text{LM}}(\theta)=-\sum_{i=1}^{L}\log P(x_{i}\mid x_{1:i})$
@@ 4 Experimental Study
In the First Stage, we train the feature maps $\phi_{q}$, $\phi_{k}$, and the gating parameter matrix $W_{\gamma}$ jointly to approximate full softmax attention and RoPE patterns. For the sliding window attention module, we use a small window size of $w=128$ and $m=4$ meta tokens. In the Second Stage, we employ Low-Rank Adaptation (LoRA) (Hu et al., 2022) for parameter-efficient fine-tuning. Specifically, LoRA is applied to the projection matrices $W_{Q}$, $W_{K}$, and $W_{V}$, with a default rank $r=8$ and scaling factor $\alpha=16$. Both stages are trained for 2 epochs, corresponding to 20M tokens per stage. We use the AdamW optimizer and a cosine learning rate schedule with a 10% linear warmup. Training is performed using Fully Sharded Data Parallelism (FSDP-2) (Zhao et al., 2023) across 8$\times$×A100 80GB GPUs. The peak learning rate is set to $1\times 10^{-3}$ for the first stage and $5\times 10^{-4}$ for the second stage. We adopt a global batch size of 8, with each example having a maximum sequence length of 2048 tokens.
@@ 4.1 Language Modeling Benchmarks
Model Training PiQA ARC-e ARC-c Hella. Wino. MMLU Avg. Avg.
Tokens (B) acc acc acc_norm acc_norm acc (5-shot) (no MMLU)
LLaMA-3-8B 15000 79.9 80.1 53.3 73.1 79.1 66.6 73.1 72.0
Hybrid Softmax
StripedHyena-Nous-7B 78.8 77.2 40.0 66.4 76.4 26.0 67.8 60.8
Zamba-7B 1000 81.4 74.5 46.6 76.4 80.2 57.7 71.8 69.5
Linearized (Keep 50% Full Attn.)
Mamba2-LLaMA-3 20 81.5 78.8 58.2 71.5 79.5 56.7 73.9 71.0
LLaMA-3-8B-Lizard (Ours) 0.04 82.2 83.1 55.9 73.6 81.4 65.1 75.2 73.5
We compare Lizard against two groups of baselines. The first group, presented in Table 1, consists of subquadratic LLMs, including models pre-trained from scratch with linear or subquadratic attention mechanisms, such as Mamba (Gu and Dao, 2023), RWKV-6 (Peng et al., 2023a), TransNormerLLM-7B (Qin et al., ), Hawk and Griffin (De et al., ), as well as linearized variants such as SUPRA (Mercat et al., 2024), LoLCATs (Zhang et al., 2024), and Mamba2LLaMA (Wang et al., 2024). Lizard consistently outperforms prior approaches, particularly on the 5-shot MMLU benchmark, where it achieves an 18% improvement over previous methods with similar extrapolation capabilities. Compared to LoLCATs, which does not generalize beyond training context, Lizard scores 8 points higher on the 5-shot MMLU.
4.2 Recall Evaluations
Figure 3: Example from the synthetic passkey retrieval dataset.
We evaluate Lizard and compare its performance against the teacher model and LoLCATs (Zhang et al., 2024), a recent state-of-the-art linearization method. Figure  4 reports the results of all three models on the associative recall test set. We find that Lizard significantly outperforms LoLCATs in both associative recall accuracy and length generalization. Notably, Lizard is able to perfectly retrieve the correct passkey across various context lengths, while LoLCATs fails when the sequence length exceeds the training window. This highlights the strength of the gated recurrent structures, which effectively compresses global context and does not rely solely on the expressiveness of local sliding window attention.
Figure 5: Throughput and memory comparison.
4.3 Generation Efficiency
@@
GLA (Yang et al., 2023) $\Gamma_{i}=\sigma\left(\mathbf{x}_{i}W_{\gamma_{1}}W_{\gamma_{2}}\right)$ $W_{\gamma_{1}}\in\mathbb{R}^{d\times 16},\;W_{\gamma_{2}}\in\mathbb{R}^{16\times d}$ 53.5
@@ 4.4 Ablation and Architectural Analysis
Effect of window and meta memory size Table 4 presents an ablation study evaluating the impact of varying the local attention window size ($w$) and the number of meta tokens ($m$) on 5-shot MMLU performance. While increasing the window size can improve performance, it does not guarantee consistent gains. For instance, performance peaks at $w=128$ for $m=4$, but drops notably at $w=256$. This indicates that simply enlarging the local attention span is not always beneficial.
5 Conclusion
We present Lizard, a linearization framework that transforms pretrained Transformers into scalable, subquadratic architectures. Our method achieves near-parity with the teacher model on standard language modeling benchmarks, including 5-shot MMLU, while enabling constant-memory inference and strong length generalization. Through extensive experiments and ablations, we demonstrate that lightweight gating designs and hybrid attention improve both performance and efficiency, making Lizard a practical solution for long-context language modeling.
Appendix A Limitations
@@
Appendix B Inference Efficiency
@@ Appendix C Ablation Study
Appendix D Experimental Details
Table 7: Hyperparameters for the experiments.
@@ Appendix D Experimental Details
Hyperparameters All model and training hyperparameters are summarized in Table 7. For the learning rate, we performed an initial sweep over {1e-2, 5e-3, 1e-3, 5e-4, 1e-4}. We did not tune the batch size. For the other designs, we adopted the default values used by prior work [Zhang et al., 2024].
Appendix E Evaluation on small-size LLMs
@@ Appendix F Sample Generations
Prompt
Why did we linearize the chungus among us?
LoLCATs Response [Zhang et al., 2024] Linearizing the chungus among us is a process of organizing and arranging the information in a logical and coherent manner. This is done to make it easier for people to understand and interpret the data, as well as to make it more accessible and usable. By linearizing the data, we are able to extract meaningful insights and patterns from the information, which can then be used to make informed decisions and take action.

### v3

@@ 3.1 First Stage: Approximating Softmax Attention for Unbounded Context
where $N$ is the number of attention layers in the model. Overall, Lizard achieves an $O\bigl(L(w+m)d+Ld^{2}\bigr)$ time and space complexity. For generation, Lizard requires only $O((w+m)^{2}d+d^{2})$ time and space for every token.
3.2 Second Stage: Aligning with Language Modeling
While the first stage ensures a high-fidelity architectural approximation, the second stage aligns the model with the downstream language modeling task. In this stage, the original softmax attention layers are replaced by Lizard attention layers, and the entire model is fine-tuned using the standard autoregressive language modeling objective: $\mathcal{L}_{\text{LM}}(\theta)=-\sum_{i=1}^{L}\log P(x_{i}\mid x_{<i})$ This step bridges the gap between structural mimicry and optimal task performance, adapting the linearized model to its end-to-end objective.
4 Hardware-Aware Algorithm for Efficient Training
To achieve maximum efficiency, Gated Linear Attention (GLA) must be computed in a parallel form on hardware accelerators like GPUs. However, the standard parallel formulation suffers from a critical numerical instability that prevents the use of low-precision formats, thereby creating a performance bottleneck. The parallel form of Gated Linear Attention is expressed as:
$\hat{\mathbf{Y}}_{\text{gate}}=\left(\left(\left(\phi(\mathbf{Q})\odot\mathbf{C}\right)\left(\frac{\phi(\mathbf{K})}{\mathbf{C}}\right)^{\top}\right)\odot\mathbf{M}\right)\mathbf{V}$
where $\mathbf{C}$ is the matrix of cumulative gate products, with each row $\mathbf{c}_{t}=\prod_{j=1}^{t}\mathbf{\Gamma}_{j}$.
The matrix form of gated linear attention is not numerically stable, as the cumulative product of gating values $\mathbf{c}_{t}$ can become extremely small, leading to underflow and instability during training for the low precision format such as bfloat16. This forces a fallback to full-precision (float32) operations, which is $2-3\times$ slower and memory intensive, preventing the use of hardware accelerators like Tensor Cores that are optimized for low-precision arithmetic.
We leverage the strictly non-negative property of the Hedgehog feature map (Zhang et al., ), $\phi(\mathbf{x})=[\exp(\mathbf{x}\mathbf{W})\oplus\exp(-\mathbf{x}\mathbf{W})]$. This exponential-based structure is critical, as it permits a stable reparameterization of the attention computation in log-space. We absorb the cumulative gate term $\mathbf{C}$ directly into the query and key projections, resulting in the following hardware-efficient formulation:
$\widetilde{\mathbf{Q}}$ $=\left[\exp(\mathbf{Q}\mathbf{W}+\log\mathbf{C})\oplus\exp(-\mathbf{Q}\mathbf{W}+\log\mathbf{C})\right]$
$\widetilde{\mathbf{K}}$ $=\left[\exp(\mathbf{K}\mathbf{W}-\log\mathbf{C})\oplus\exp(-\mathbf{K}\mathbf{W}-\log\mathbf{C})\right]$
By shifting the unstable gating contributions into the feature space, this approach transforms the core operation into a standard General Matrix Multiplication (GEMM): $\hat{\mathbf{Y}}_{\text{gate}}=\left(\left(\widetilde{\mathbf{Q}}\widetilde{\mathbf{K}}^{\top}\right)\odot\mathbf{M}\right)\mathbf{V}$ which aligns with the native mma.sync Tensor Core instruction. This avoids custom CUDA kernels and leverages highly optimized GEMM routines in cuBLAS and FlashAttention backends (Dao, 2023), avoids a full-precision fallback. As a result, it improves training throughput by up to 32% compared to prior gated attention implementations (Yang et al., 2024)  [^1] https://github.com/fla-org/flash-linear-attention.
@@ 4 Hardware-Aware Algorithm for Efficient Training
Model Training PiQA ARC-e ARC-c Hella. Wino. MMLU Avg. Avg.
Tokens (B) acc acc acc_norm acc_norm acc (5-shot) (no MMLU)
LLaMA-3-8B 15000 79.9 80.1 53.3 73.1 79.1 66.6 73.1 72.0
Hybrid Softmax
StripedHyena-Nous-7B 78.8 77.2 40.0 66.4 76.4 26.0 67.8 60.8
Zamba-7B 1000 81.4 74.5 46.6 76.4 80.2 57.7 71.8 69.5
Linearized (Keep 50% Full Attn.)
Mamba2-LLaMA-3 20 81.5 78.8 58.2 71.5 79.5 56.7 73.9 71.0
LLaMA-3-8B-Lizard (Ours) 0.04 82.2 83.1 55.9 73.6 81.4 65.1 75.2 73.5
@@ 5 Experimental Study
[…]
@@ 5.1 Language Modeling Benchmarks
GLA (Yang et al., 2024) $\Gamma_{i}=\sigma\left(\mathbf{x}_{i}W_{\gamma_{1}}W_{\gamma_{2}}\right)$ $W_{\gamma_{1}}\in\mathbb{R}^{d\times 16},\;W_{\gamma_{2}}\in\mathbb{R}^{16\times d}$ 53.5
@@
We compare Lizard against two groups of baselines. The first group, presented in Table 1, consists of subquadratic LLMs, including models pre-trained from scratch with linear or subquadratic attention mechanisms, such as Mamba (Gu and Dao, 2023), RWKV-6 (Peng et al., 2023a), TransNormerLLM-7B (Qin et al., ), Hawk and Griffin (De et al., ), as well as linearized variants such as SUPRA (Mercat et al., 2024), Mamba2LLaMA (Wang et al., 2024), LoLCATs (Zhang et al., 2025), Liger (Lan et al., 2025). Lizard consistently outperforms prior approaches, particularly on the 5-shot MMLU benchmark, where it achieves an 18% improvement over previous methods with similar extrapolation capabilities. Compared to LoLCATs and Liger, which does not generalize beyond training context, Lizard scores 9.4 and 24.5 points higher on the 5-shot MMLU, respectively.
Figure 3: Example from the synthetic passkey retrieval dataset.
5.2 Recall Evaluations
We evaluate Lizard and compare its performance against the teacher model and LoLCATs (Zhang et al., 2025), a recent state-of-the-art linearization method. Figure  2 reports the results of all three models on the associative recall test set. We find that Lizard significantly outperforms LoLCATs in both associative recall accuracy and length generalization. Notably, Lizard is able to perfectly retrieve the correct passkey across various context lengths, while LoLCATs fails when the sequence length exceeds the training window. This highlights the strength of the gated recurrent structures, which effectively compresses global context and does not rely solely on the expressiveness of local sliding window attention.
5.3 Generation Efficiency
Figure 4: Throughput and memory comparison.
@@ 5.4 Ablation and Architectural Analysis
Effect of window and meta memory size Table 4 presents an ablation study evaluating the impact of varying the local attention window size ($w$) and the number of meta tokens ($m$) on 5-shot MMLU performance. While increasing the window size can improve performance, it does not guarantee consistent gains. For instance, performance peaks at $w=128$ for $m=4$, but drops notably at $w=256$. This indicates that simply enlarging the local attention span is not always beneficial. This suggests that an overly large local window may introduce noise or distract from the compressed global context provided by the gated attention module, highlighting a delicate trade-off between local precision and global context.
6 Limitations
@@
Appendix A Inference Efficiency
@@ Appendix B Ablation Study
Table 6: Effect of LoRA Rank.
Appendix C Experimental Details
@@
Table 7: Hyperparameters for the experiments.
Hyperparameters All model and training hyperparameters are summarized in Table 7. For the learning rate, we performed an initial sweep over {1e-2, 5e-3, 1e-3, 5e-4, 1e-4}. We did not tune the batch size. For the other designs, we adopted the default values used by prior work (Zhang et al., 2025).
Appendix D Evaluation on small-size LLMs
Evaluation on small-size LLMs is shown in Table 8.
@@ Appendix D Evaluation on small-size LLMs
Appendix E Evaluation on LongBench
The results below show that Lizard consistently achieves better or comparable performance to the original teacher model on representative LongBench tasks as shown in Table 9
Dataset Llama-3.1-8B -Instruct Llama-3.1-8B -Instruct-Lizard
2WikiMultihopQA 16.8 16.2
HotpotQA 18.3 19.6
MultiNews 28.5 27.8
PassageRetrieval-en 98.1 97.9
LCC 51.3 53.6
RepoBench-P 48.6 52.7
Table 9: Comparison of Llama-3.1-8B-Instruct and its Lizard variant on various benchmarks.
@@ Appendix F Sample Generations
Prompt Why did we linearize the chungus among us?
LoLCATs Response (Zhang et al., 2025) Linearizing the chungus among us is a process of organizing and arranging the information in a logical and coherent manner. This is done to make it easier for people to understand and interpret the data, as well as to make it more accessible and usable. By linearizing the data, we are able to extract meaningful insights and patterns from the information, which can then be used to make informed decisions and take action.